# Dispensing with optimal control: a new approach for the pricing and management of share buyback contracts

### v2

@@ Abstract
This paper introduces a novel methodology for the pricing and management of share buyback contracts, overcoming the limitations of traditional optimal control methods, which frequently encounter difficulties with high-dimensional state spaces and the intricacies of selecting appropriate risk penalty or risk aversion parameter. Our methodology applies optimized heuristic strategies to maximize the contract’s value. The computation of this value utilizes classical methods typically used for pricing path-dependent Bermudan options. Additionally, our approach naturally leads to the formulation of a hedging strategy.
@@ Introduction
Over the past decades, share buybacks have globally surged, even outpacing dividend increases, though they faced a downturn in 2023 in the US due to rising interest rates, as US companies frequently finance buybacks with borrowed funds. The shift from fixed-price tender offers and Dutch auctions to predominantly open market repurchases (OMRs) in the 90s, followed by the rise of Accelerated Share Repurchase (ASR) contracts and buyback mandates, historically marks a significant evolution in how buybacks are conducted. ASR setups are more common in the US. This approach involves a company rapidly repurchasing its own shares through a contract with an investment bank. Typically, the bank initially borrows the shares to deliver them to the company, then gradually closes its short position by buying shares in the open market. In contrast, the mandate format, which is more prevalent in Europe, allows the bank to slowly purchase shares on the market and progressively deliver them to the company.
@@
In both ASRs and mandate share buyback contracts, the bank’s compensation is often tied to a price benchmark, typically the Volume Weighted Average Price (VWAP) over a period decided dynamically by the bank which faces an optimal stopping time problem. Competitive pricing strategies in these contracts hinge on offering discounts relative to this benchmark. A company may request specific features in these contracts, and banks, particularly those advising the firm, can propose complex terms. Recent years have seen an increase in original features within contracts such as floating notional, lookback features, day exclusion upon insufficient daily volume or high price increase and subsequent maturity extensions or notional reduction. This trend emphasizes the importance of quantitative analysts in accurately pricing and managing these contracts.
@@
Often seen as variants of optimal execution problems, the pricing and management of share buyback contracts have mainly been addressed with the tools of stochastic optimal control. Jaimungal and his coauthors in [7], followed by Guéant and his coauthors in [5], and then Guéant in [3] have indeed proposed several models relying on optimal control tools for both fixed quantity and fixed notional ASR contracts. These models ultimately involve solving Hamilton-Jacobi-Bellman equations, or simply Bellman equations, using grid or sometimes tree methods. While these methods are efficient in simple cases and have been adopted by several banks, they encounter three significant challenges. First, dealing with complex contracts or advanced price dynamics necessitates a high-dimensional state space, with mandates posing greater difficulties in this regard than ASR contracts. Although neural networks can potentially overcome the curse of dimensionality (see [4] and [6] for techniques inspired by reinforcement learning ideas), their opaque decision-making process and unpredictable behavior with extreme state values often lead to their rejection by practitioners. The second issue concerns the selection of an appropriate risk penalty and/or risk aversion parameter to mitigate contract execution risks. This choice significantly influences both optimal strategies and pricing. Moreover, the concept of pricing tied to optimal control methods relates to indifference pricing, seldom used by practitioners. Third, and related to the previous points, optimal control tools often conflict with the pricing and hedging frameworks that are commonly found in the libraries of most investment banks.
@@
To address these limitations, Baldacci et al. introduced in [2] an alternative approach that simplifies the problem through a heuristic repurchase strategy. This strategy is designed to intuitively align with the expected directionality of parameter effects, reducing the complexity to the more manageable task of pricing exotic American or Bermudan options using conventional methods. Our research aims to expand upon this foundation, focusing on refining and optimizing heuristic strategies, thus offering a more detailed insight into the management of share buyback strategies and contracts.
@@
In our first and simple framework, we consider the case of a theoretical buyback contract with notional $F>0$, maturity $T>0$, and no constraint whatsoever. We assume that there is no friction in the market, i.e. the trader can buy or sell any number of shares at any time without transaction costs or market impact. [^4] We diverge here from the traditional perspective commonly found in the academic literature, which treats buyback issues primarily as execution challenges. This departure is partly based on the idea that the market has already assimilated the informational content of trades. Moreover, we posit that the option component inherent in buyback contracts holds more significance than the aspects related to execution costs.
@@
The Longstaff-Schwartz method is a Monte-Carlo method based on the above dynamic programming principle. One first needs to simulate $M\in\mathbb{N}^{*}$ sample paths of the process $\left(Y^{1}_{t_{n}}\right)_{n},\ldots,\left(Y^{M}_{t_{n}}\right)_{n}$. For each $m\in\{1,\ldots,M\}$, we have the following terminal condition:
@@
:=${\cal F}:=\left\{f(\cdot,\cdot;\beta):[0,T]\times\mathbb{R}_{+}^{*}\rightarrow\mathbb{R}\left|\beta\subset B\right.\right\},$
where $B\subset\mathbb{R}^{d}.$ We then define for each $\beta\in B$ the stopping time
@@
with the convention $\inf\emptyset=T$. We can approximate the value of $\mathbb{E}^{\mathbb{Q}}\left[Y_{\tau^{\beta}}\right]$ with a standard Monte-Carlo method, and then use a grid-search algorithm in order to optimize over $\beta\in B$. If the family ${\cal F}$ is rich enough, the approximate problem
@@
In this part, we compare numerically the performance of the strategies obtained with the three methods described above, respectively. We consider the following parameters:
@@
:=${\cal F}:=\left\{f(\cdot,\cdot;\beta):[0,T]\times\mathbb{R}_{+}^{*}\rightarrow\mathbb{R}\left|\beta\subset\mathbb{R}^{3}\right.\right\},$
@@
We then test the strategies obtained with the different methods using a Monte-Carlo simulation with $10,000$ trajectories. The average payoff and standard deviation for each method is reported in Table 1. The distributions of the values obtained for the different methods are plotted in Figures 1, 2 and 3, respectively. It is remarkable that the three strategies yield very close results both in terms of expected value and standard deviation. The distributions under $\mathbb{Q}$ of the resulting payoffs are spread below $1$ and have a spike above $1$. In all cases, the variability of the payoffs can be hedged away using standard mathematical finance techniques.
@@
In what follows we consider a buyback contract with a floating maturity comprised between $N_{min}>0$ and $N_{max}\geq N_{min}$, and a floating notional comprised between $F_{min}>0$ and $F_{max}\geq F_{min}$. The trader can choose to exercise early at any time after a given date $N_{ex}$, with $0\leq N_{ex}\leq N_{min}$.
@@
There may be a cap $S_{max}$ on the price such that, if $S_{n}>S_{max}$, the day is suspended, i.e. excluded from the computation of the average price. [^5] A barrier $V_{min}$ might also be enforced such that, if the total market volume $V_{n}$ on day $n$ is such that $V_{n}<V_{min}$, the day $n$ is suspended. However, this case is not treated in this paper, as the variability of $V_{n}$ is not modeled and can hardly be hedged. For all $n$, let us denote by $C_{n}$ the number of suspended days up to time $n$. Initially, the maturity of the contract is given by $N_{min}$. When a trading day is suspended, the maturity of the program is extended by one day up to the maximal maturity $N_{max}$. If the number of suspended days exceeds $N_{max}-N_{min}$, $F_{min}$ is reduced pro-rata by the excess number of suspended days, i.e.
@@
Therefore, the trader just solves the following problem
@@
These two strategies are then optimized over $\alpha,\beta,\gamma$, and $a$ using the optuna library, and their respective performance are reported in the table below (corresponding to 2,000 simulations). We see in particular that the two strategies clearly outperform the benchmarks in terms of average PnL. We also observe that adding more parameters to the early execution strategy does not improve the performance.
@@
    $v_{n}=\frac{2}{1+\exp\left\{a+a_{1}\frac{T_{n}-n}{T_{n}}\right\}}\times\frac{\bar{F}-X_{n}}{S_{n}(T_{n}-n)},$
@@
    $v_{n}=\frac{2}{1+\exp\left\{a+a_{1}\frac{T_{n}-n}{T_{n}}+a_{2}\frac{\max(N_{ex}-n,0)}{N_{ex}}\right\}}\times\frac{\bar{F}-X_{n}}{S_{n}(T_{n}-n)},$
@@
    $v_{n}=\frac{2}{1+\exp\left\{a+a_{1}\frac{T_{n}-n}{T_{n}}+b\left(\frac{A}{S}-1\right)\right\}}\times\frac{\bar{F}-X_{n}}{S_{n}(T_{n}-n)},$
@@
    $v_{n}=\frac{2}{1+\exp\left\{a+a_{1}\frac{T_{n}-n}{T_{n}}+\left(b+b_{1}\frac{T_{n}-n}{T_{n}}\right)\left(\frac{A}{S}-1\right)+c\frac{F_{min}-X_{n}}{S_{n}v_{max}(T_{n}-n)}\right\}}\times\frac{\bar{F}-X_{n}}{S_{n}(T_{n}-n)},$
@@
In conclusion, our research marks a significant shift from traditional approaches that relied on stochastic optimal control tools for pricing and managing share buyback contracts, which were originally inspired by optimal execution methods. By refining and optimizing heuristic strategies, we successfully address the limitations posed by high-dimensional state spaces and the complex selection of risk parameters. Our method simplifies the issue to pricing and hedging a Bermudan payoff, making it tractable with conventional financial tools. This approach not only builds a bridge between theory and application in financial markets but also separates the strategies for delivering shares to the company from the hedging of the payoff, potentially elucidating execution patterns observed in the market.
@@
This research has been conducted with the support of the Research Initiative “Modélisation des marchés actions, obligations et dérivés” financed by HSBC France under the aegis of the Europlace Institute of Finance. Manuel Abellan-Lopez (HSBC), Jérôme Lemue (HSBC) and Stéphane Quentin (HSBC) also deserve special thanks for the discussions we had with them. The views expressed are those of the authors and do not necessarily reflect the views or the practices at HSBC.

### v3

@@ Abstract
This paper introduces a novel methodology for the pricing and management of share buyback contracts, overcoming the limitations of traditional optimal control methods, which frequently encounter difficulties with high-dimensional state spaces and the intricacies of selecting appropriate risk penalty or risk aversion parameter. Our methodology applies optimized heuristic strategies to maximize the contract’s value. The computation of this value utilizes classical methods typically used for pricing path-dependent options. Additionally, our approach naturally leads to the formulation of a $\Delta$-hedging strategy and disentangles therefore the repurchase strategy from the hedging of the payoff.
@@ Introduction
Over the past decades, share buybacks have globally surged, even outpacing dividend increases, though they faced a downturn in 2023 in the US due to rising interest rates, as US companies frequently finance buybacks with borrowed funds. The shift from fixed-price tender offers and Dutch auctions to predominantly open market repurchases (OMRs) in the 90s, followed by the rise of Accelerated Share Repurchase (ASR) contracts and buyback mandates, historically marks a significant evolution in how buybacks are conducted. ASR setups are more common in the US. This approach involves a company rapidly repurchasing its own shares through a contract with an investment bank. Typically, the bank initially borrows the shares to deliver them to the company and then gradually closes its short position by buying shares in the open market. In contrast, the mandate format, which is more prevalent in Europe, allows the bank to slowly purchase shares on the market and progressively deliver them to the company.
@@
In both ASR contracts and share buyback mandates, the bank’s compensation is often tied to a price benchmark, typically the Volume Weighted Average Price (VWAP) over a period decided dynamically by the bank itself, which thus faces an optimal stopping time problem. Competitive pricing strategies in these contracts hinge on offering discounts relative to this benchmark. A company may request specific features in these contracts, and banks, particularly those advising the firm, can propose complex terms. Recent years have seen an increase in original features within contracts such as floating notional, lookback features, day exclusion upon insufficient daily volume or high price increase and subsequent maturity extensions or notional reduction. This trend emphasizes the importance of quantitative analysts in accurately pricing and managing these contracts.
@@
Often seen as variants of optimal execution problems, the pricing and management of share buyback contracts have mainly been addressed with the tools of stochastic optimal control. Jaimungal and his coauthors in [7], followed by Guéant and his coauthors in [5], and then Guéant in [3] have indeed proposed several models relying on optimal control tools for both fixed quantity and fixed notional ASR contracts. These models ultimately involve solving Hamilton-Jacobi-Bellman equations, or simply Bellman equations, using grid or sometimes tree methods. While these methods are efficient in simple cases and have been adopted by several banks, they encounter three significant challenges. First, dealing with complex contracts or advanced price dynamics necessitates a high-dimensional state space, with mandates posing greater difficulties in this regard than ASR contracts. Although neural networks can potentially overcome the curse of dimensionality (see [4] and [6] for techniques inspired by reinforcement learning ideas), their opaque decision-making process and unpredictable behavior with extreme state values often lead to their rejection by practitioners. The second issue concerns the selection of an appropriate risk penalty and/or risk aversion parameter to mitigate contract execution risks. This choice significantly influences both optimal strategies and pricing. Moreover, the concept of pricing tied to optimal control methods is indifference pricing, a concept seldom used by practitioners. Third, and related to the previous points, optimal control tools often conflict with the pricing and hedging frameworks that are commonly found in the libraries of most investment banks.
@@
To address these limitations, Baldacci et al. introduced an alternative approach in [2] that simplifies the problem through a heuristic repurchase strategy. This strategy intuitively aligns with the expected directionality of parameter effects, reducing the complexity to the more manageable task of pricing a path-dependent option with American or Bermudean features. Our research builds upon this foundation by refining and optimizing heuristic strategies, thereby providing more detailed insights into the management of share buyback contracts. Specifically, we maximize the expected value under $\mathbb{Q}$ of the payoff associated with heuristic repurchase and stopping strategies using state-of-the-art hyperparameter optimization tools. The resulting payoff can then be hedged dynamically using standard $\Delta$-hedging techniques available in conventional pricing libraries. Notably, our approach disentangles the repurchase strategy from the $\Delta$ of the payoff.
@@
In our first and simple framework, we consider the case of a theoretical buyback contract with notional $F>0$, maturity $T>0$, and no constraint whatsoever. We assume that there is no friction in the market, i.e. the trader can buy or sell any number of shares at any time without transaction costs or market impact. [^4] We diverge here from the traditional perspective commonly found in the academic literature, which treats buyback issues primarily as execution challenges. This departure is partly based on the idea that the market has already assimilated the informational content of trades. Moreover, we posit that the option component inherent in buyback contracts holds more significance for most assets than the aspects related to execution costs.
@@
The Longstaff-Schwartz method is a Monte Carlo method based on the above dynamic programming principle. One first needs to simulate $M\in\mathbb{N}^{*}$ sample paths of the process $\left(Y^{1}_{t_{n}}\right)_{n},\ldots,\left(Y^{M}_{t_{n}}\right)_{n}$. For each $m\in\{1,\ldots,M\}$, we have the following terminal condition:
@@
:=${\cal F}:=\left\{f(\cdot,\cdot;\beta):[0,T]\times\mathbb{R}_{+}^{*}\rightarrow\mathbb{R}\left|\beta\in B\right.\right\},$
where $B\subset\mathbb{R}^{d}.$ For each $\beta\in B$, we define the stopping time
@@
with the usual convention $\inf\emptyset=T$. We can then approximate the value of $\mathbb{E}^{\mathbb{Q}}\left[Y_{\tau^{\beta}}\right]$ with a standard Monte Carlo method, and then use state-of-the-art algorithms [^5] In the numerical examples below we used the Quasi-Monte Carlo (QMC) and the Tree-structured Parzen Estimator (TPE) samplers of the optuna python library. in order to optimize that value over $\beta\in B$. If the family ${\cal F}$ is rich enough, the approximate problem
@@
In this part, we compare numerically the performance of the strategies obtained with the three methods described above. We consider the following parameters:
@@
:=${\cal F}:=\left\{f(\cdot,\cdot;\beta):[0,T]\times\mathbb{R}_{+}^{*}\rightarrow\mathbb{R}\left|\beta\in\mathbb{R}^{3}\right.\right\},$
@@
We then test the strategies obtained using the three different methods through a Monte Carlo simulation with 10,000 trajectories. The average payoff and standard deviation under $\mathbb{Q}$ for each method are reported in Table 1. Remarkably, the three strategies yield very similar results in terms of both expected value and standard deviation. The distributions under $\mathbb{Q}$ of the payoffs resulting from the different methods are plotted in Figures 1, 2, and 3. The portion of the distributions below 1 is quite spread out, while a spike is observed above 1 in the three cases. Naturally, the risk of the payoff can be hedged away using standard $\Delta$-hedging, and that is why we focused on the maximization of the expected value under $\mathbb{Q}$.
@@
In what follows we consider a buyback contract with a floating maturity comprised between $N_{min}>0$ and $N_{max}\geq N_{min}$, and a floating notional comprised between $F_{min}>0$ and $F_{max}\geq F_{min}$. The trader can choose to exercise early after a given date $N_{ex}$, with $0\leq N_{ex}\leq N_{min}$.
@@
There may be a cap $S_{max}$ on the price such that, if $S_{n}>S_{max}$, the day is suspended, i.e. excluded from the computation of the average price. [^6] A barrier $V_{min}$ might also be enforced: if the total market volume $V_{n}$ on day $n$ is such that $V_{n}<V_{min}$, the day $n$ is suspended. However, this case is not treated in this paper, because the variability of $V_{n}$ is not modeled and can hardly be hedged. For all $n$, let us denote by $C_{n}$ the number of suspended days up to time $n$. Initially, the maturity of the contract is given by $N_{min}$. When a trading day is suspended, the maturity of the program is extended by one day up to the maximal maturity $N_{max}$. If the number of suspended days exceeds $N_{max}-N_{min}$, $F_{min}$ is reduced pro-rata by the excess number of suspended days, i.e.
@@
Therefore, the trader just solves the following problem [^9] Once again, we maximize the expected value under $\mathbb{Q}$ because we implicitly assume that the resulting payoff will be $\Delta$-hedged.
@@
These two strategies are then optimized over $\alpha,\beta,\gamma$, and $a$ using the TPE sampler of the optuna python library and their respective performance are reported in the table below corresponding to a Monte Carlo simulation with 2,000 trajectories. [^10] Heuristic strategies are optimised upon at the outset and then followed deterministically when we carry out Monte Carlo simulations. We see in particular that the two strategies clearly outperform the benchmarks in terms of average PnL. We also observe that adding more parameters to the early execution strategy does not improve the performance.
@@
    $v_{n}=\frac{2}{1+\exp\left(a+a_{1}\frac{T_{n}-n}{T_{n}}\right)}\times\frac{\bar{F}-X_{n}}{S_{n}(T_{n}-n)},$
@@
    $v_{n}=\frac{2}{1+\exp\left(a+a_{1}\frac{T_{n}-n}{T_{n}}+a_{2}\frac{\max(N_{ex}-n,0)}{N_{ex}}\right)}\times\frac{\bar{F}-X_{n}}{S_{n}(T_{n}-n)},$
@@
    $v_{n}=\frac{2}{1+\exp\left(a+a_{1}\frac{T_{n}-n}{T_{n}}+b\left(\frac{A}{S}-1\right)\right)}\times\frac{\bar{F}-X_{n}}{S_{n}(T_{n}-n)},$
@@
    $v_{n}=\frac{2}{1+\exp\left(a+a_{1}\frac{T_{n}-n}{T_{n}}+\left(b+b_{1}\frac{T_{n}-n}{T_{n}}\right)\left(\frac{A}{S}-1\right)+c\frac{F_{min}-X_{n}}{S_{n}v_{max}(T_{n}-n)}\right)}\times\frac{\bar{F}-X_{n}}{S_{n}(T_{n}-n)},$
@@
In conclusion, our research marks a significant shift from traditional approaches that relied on stochastic optimal control tools for pricing and managing share buyback contracts, which were originally inspired by optimal execution methods. By refining and optimizing heuristic strategies, we successfully address the limitations posed by high-dimensional state spaces and the complex selection of risk parameters. Our method simplifies the issue to pricing and hedging a path-dependent payoff, making it tractable with conventional financial tools. This approach not only builds a bridge between theory and application in financial markets but also separates the strategies for delivering shares to the company from the hedging of the payoff, potentially elucidating execution patterns observed in the market.
@@
This research has been conducted with the support of the Research Initiative “Modélisation des marchés actions, obligations et dérivés” financed by HSBC France under the aegis of the Europlace Institute of Finance. Manuel Abellan-Lopez (HSBC), Jérôme Lemue (HSBC) and Stéphane Quentin (HSBC) deserve special thanks for the discussions we had with them. The views expressed are those of the authors and do not necessarily reflect the views or the practices at HSBC.